OpenStudy (anonymous):

inverse of f(x)=x^3+1?

Parth (parthkohli):

You do the exact opposite thing. In this case, the inverse of adding is subtracting and the inverse of cubing is taking the cube root. So, we can say: \(\Large \color{midnightblue}{\rightarrow f^{-1}(x) = \sqrt[3]{x} - 1}\)

OpenStudy (anonymous):

oops

Parth (parthkohli):

Did I do it wrong?

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

you need to do the opposite thing in the reverse order

Parth (parthkohli):

Oops...

OpenStudy (anonymous):

your function says 1) cube 2) add one inverse says 1) subtract 1 2) take the cubed root \[f^{-1}(x)=\sqrt[3]{x-1}\]

Parth (parthkohli):

Okay!

Parth (parthkohli):

I got it, thank you @satellite73

OpenStudy (anonymous):

if you like, write \[y=x^3+1\] the switch x and y get \[x=y^3+1\] solve for y via \[x-1=y^3\] \[y=\sqrt[3]{x-1}\]

OpenStudy (anonymous):

yw